A passenger train's speed is 60 mi/h, and a freight train's speed is 40 mi/h. The passenger train travels the same distance in 1.5 h less time than the freight train. How long does each train take to make the trip?
Answer by Paul(988)

A passenger train's speed is 60 mi/h, and a freight train's speed is 40 mi/h. The passenger train travels the same distance in 1.5 h less time than the freight train. How long does each train take to make the trip?
60(t-1.5)=40t
60t-90=40t
-90=-20t
t=4.5
Therefore one takes 4.5h and the other takes 3h. Sorry I do not have time to tell you which one is which.
The distance they both cover is 180miles